OpenStudy (anonymous):

what is the factored forms of the following expressions 1. d^2+12d+32 2. d^2-12d-32 3. x^2-x-42 4. d^2-18d+80 5.x^2+3xy-4y^2

OpenStudy (lieutenantgeneral):

1.) d= -8 d = -4 2.) d = 6 - 2square root 17 d = 2(3 + square root 17 3.) x = -6 x = 7 4.)d = 8 d = 10 5.)x = 4n y = -n

OpenStudy (anonymous):

thank you

OpenStudy (lieutenantgeneral):

you welcome